For the real valued functions f and g are defined below find fog and gof Where

f(x) ="\sqrt{x+1}" and g(x)= "\frac{1}{x^2}"

1
Expert's answer
2020-11-03T18:18:42-0500

fog=f(g(x))=f(1/x2)="\sqrt{1/x^2+1}" ="\cfrac{\sqrt{x^2+1}}{\vert x\vert} ; x\isin\R"

gof=g(f(x))="g(\sqrt{x+1})" = "\cfrac{1}{x+1} ; x>-1"